# Comparison of Extended Lubrication Theories for Stokes Flow

### v1

@@ Abstract
Lubrication theory leverages the assumption of a long and thin fluid domain to formulate a linearized approximation to the Navier-Stokes equations. Models in extended lubrication theory consider relaxations of the thin film assumption, leading to the inclusion of higher order terms. However, such models are sensitive to large surface gradients which lead the assumptions of the model to break down. In this paper, we present a formulation of extended lubrication theory, and compare our models with several existing models, along with the numerical solution to the Stokes equations. The error in pressure and velocity is characterized for a variety of fluid domain geometries. Our results indicate that the new solution is suitable for a wide range of geometries. Nonetheless, the magnitude of surface variation and the length scale ratio are both important factors influencing the accuracy of the extended lubrication theory models.
@@ 1 Introduction
The classical lubrication assumption of a long and thin fluid domain is used to justify neglecting the inertial velocity and cross-film pressure gradient terms in the Navier-Stokes equations. The resulting Reynolds equation is a dramatic simplification of the original governing equations. More recently, variations on the thin film assumption have given rise to higher order approximations for lubrication theory [16, 15, 14]. These extended lubrication theory models retain certain linearity properties as in classical lubrication theory, while nevertheless being applicable to a broader scope of geometries in the low Reynolds number limit.
@@
Previous research on the flows modeled by lubrication theory has highlighted the sensitivity of lubrication theory to variations of the fluid film height. In particular, in the limit of zero Reynolds number flows, discrepancies between lubrication theory and Stokes flow increase with both the magnitude and frequency of large gradients in the height [4, 2, 5]. The total pressure drop modeled by the Reynolds equation may be underestimated due to additional pressure losses at points of sudden expansion that are not accurately modeled [6, 5]. Furthermore, notable flow features of corner recirculation seen in Stokes flow are not accurately captured with the Reynolds equation [1, 2, 13].
@@
In this paper, we consider classical lubrication theory (the Reynolds equation), and more recent variations on lubrication theory which we refer to as extended lubrication theory [14] and/or perturbed lubrication theory [16, 10, 15] in the case of models based on perturbation theory. We asses the performance of these models at low Reynolds numbers by comparison to the Stokes equations. We propose a formulation of extended lubrication theory that builds on the model of Takeuchi & Gu (2019); our approach modifies the treatment of the velocity boundary conditions and ensures incompressibility. The pressure and velocity obtained from each model are compared in a variety of geometries: a logistic step, a backward facing step, and two versions of a triangular slider that we refer to as positive and negative based on the texture to be defined later on. Through these examples, we assess the sensitivity of each model to varying magnitudes of surface variation and discontinuities in the surface gradient.
@@
Our results indicate that the extended and perturbed lubrication theory models provide significant improvements on classical lubrication theory in the case of small to moderate surface variation. For the logistic step, our proposed solution in extended lubrication theory has the smallest error in velocity for small slopes, and the smallest error in pressure at moderate slopes. For the triangular slider, the perturbed solutions have significantly reduced error in pressure, particularly in the case of negative texturing. In general, we find that for examples with a large surface gradient, the extended and perturbed solutions may over-correct the cross-film pressure gradient.
@@ 2.1 Lubrication Theory
We begin with a review of classical lubrication theory. This treatment may be found in various textbooks [8], and we include it here to establish notation and to make the presentation self-contained. The Navier-Stokes equations,
@@
serve as the governing equations for a two dimensional incompressible fluid. The pressure is $p$, the velocity is $(u,v)$, and the kinematic viscosity is $\nu=\eta/\rho$, where $\eta$ is the bulk viscosity and $\rho$ is the constant density.
Lubrication theory provides a linearized approximation to the Navier-Stokes equations valid for geometries in which the fluid is constrained to a relatively thin domain. For the two dimensional fluid domain $(x,y)\in[x_{0},x_{f}]\times[0,h(x)]$, the characteristic length scales are $L_{x}=x_{f}-x_{0}$ and $L_{y}=\max h(x)>0$. Given a prescribed constant flux $\mathcal{Q}\neq 0$, the characteristic velocities are $U_{*}=\mathcal{Q}/L_{y}$ and $V_{*}=\mathcal{Q}/L_{x}$, and the characteristic pressure is $P_{*}=\eta U_{*}L_{x}/L_{y}^{2}$. The nondimensionalized variables are $\bar{x}=x/L_{x}$, $\bar{y}=y/L_{y}$, $\bar{u}=u/U_{*}$, $\bar{v}=v/V_{*}$, and $\bar{p}=p/P_{*}$. The Reynolds number is $\text{Re}=\rho U_{*}L_{y}/\eta$ and the length scale ratio is denoted $\varepsilon=L_{y}/L_{x}$. The Navier-Stokes equations Eqs. 1 and 2 in dimensionless terms are then,
(4) $\frac{\partial\bar{p}}{\partial\bar{x}}$ $=\Big(\varepsilon^{2}\frac{\partial^{2}\bar{u}}{\partial\bar{x}^{2}}+\frac{\partial^{2}\bar{u}}{\partial\bar{y}^{2}}\Big)-\varepsilon^{2}\text{Re}\Big(\bar{u}\frac{\partial\bar{u}}{\partial\bar{x}}+\bar{v}\frac{\partial\bar{u}}{\partial\bar{y}}\Big),$
(5) $\frac{\partial\bar{p}}{\partial\bar{y}}$ $=\varepsilon^{2}\Big(\varepsilon^{2}\frac{\partial^{2}\bar{v}}{\partial\bar{x}^{2}}+\frac{\partial^{2}\bar{v}}{\partial\bar{y}^{2}}\Big)-\varepsilon^{4}\text{Re}\Big(\bar{u}\frac{\partial\bar{v}}{\partial\bar{x}}+\bar{v}\frac{\partial\bar{v}}{\partial\bar{y}}\Big),$
@@
(6) $\frac{\partial p}{\partial x}$ $=\eta\frac{\partial^{2}u}{\partial y^{2}},$
(7) $\frac{\partial p}{\partial y}$ $=0.$
Together with incompressibility Eq. 3, Eqs. 6 and 7 constitute the governing equations for lubrication theory.
(8) $u(x,0)=\mathcal{U}$ $u(x,h(x))=0,$
(9) $v(x,0)=0$ $v(x,h(x))=0,$
where $\mathcal{U}$ is the constant relative velocity between the surfaces. We prescribe a constant flux $\mathcal{Q}$, corresponding to the inlet pressure gradient, and set the outlet pressure to be zero,
(10) $\frac{dp}{dx}\bigg|_{x_{0}}=\frac{-12\eta\mathcal{Q}}{\big[h(x_{0})\big]^{3}}+\frac{6\eta\mathcal{U}}{\big[h(x_{0})\big]^{2}},$ $p(x_{L},y)=0.$
The velocity $u=u_{\text{Re}}$ is determined through integration of Eq. 6 and applying the boundary conditions Eq. 8,
(11) $u(x,y)=\frac{1}{2\eta}\frac{dp}{dx}\Big(y^{2}-h(x)y\Big)+\frac{\mathcal{U}}{h(x)}\Big(h(x)-y\Big).$
The velocity $v=v_{\text{Re}}$ is determined from incompressibility Eq. 3 and applying the boundary conditions Eq. 9,
(12) $v(x,y)=\frac{-1}{6\eta}\frac{d^{2}p}{dx^{2}}y^{3}+\frac{1}{2}\Bigg(\frac{1}{2\eta}\Big(\frac{d^{2}p}{\partial x^{2}}h(x)+\frac{dp}{dx}\frac{dh}{dx}\Big)-\frac{\mathcal{U}}{[h(x)]^{2}}\frac{dh}{dx}\Bigg)y^{2}.$
The condition of constant flux $\mathcal{Q}$,
(13) $\mathcal{Q}=\int_{0}^{h(x)}u(x,y)\,dy=\frac{-1}{12\eta}\Bigg(\Big[h(x)\Big]^{3}\frac{dp}{dx}-6\eta\mathcal{U}h(x)\Bigg),$
and the boundary condition $v(x,h)=0$ are satisfied exactly when $p(x)$ satisfies the Reynolds equation,
(14) $\frac{d}{dx}\Bigg[\Big[h(x)\Big]^{3}\frac{dp}{dx}\Bigg]=6\eta\,\mathcal{U}\frac{dh}{dx}.$
The solution of pressure $p=p_{\text{Re}}$ for Reynolds equation is determined using a second-order accurate finite difference scheme, see Appendix A.
@@ 2.2 Perturbed Lubrication Theory
The dimensionless Navier-Stokes equations Eqs. 4 and 5 motivate the approach of an $\varepsilon$-perturbed solution. Given a sufficiently small Reynolds number such that $\varepsilon^{2}\text{Re}\ll 1$, perturbed lubrication theory assumes a solution of pressure and velocity of the form,
(15) $\bar{p}(\bar{x},\bar{y})$ $=P_{0}+\varepsilon^{2}P_{2}+\varepsilon^{4}P_{4},$
(16) $\bar{u}(\bar{x},\bar{y})$ $=U_{0}+\varepsilon^{2}U_{2}+\varepsilon^{4}U_{4},$
(17) $\bar{v}(\bar{x},\bar{y})$ $=V_{0}+\varepsilon^{2}V_{2}+\varepsilon^{4}V_{4}.$
The order $\varepsilon^{k}$ solution ($\varepsilon^{k}$-PLT) assumes a long and thin domain characterized by $\varepsilon^{k}\sim\mathcal{O}(1)$ and such that higher powers of $\varepsilon$ are negligible; the order $\varepsilon^{0}$-PLT solution is the solution of lubrication theory and the Reynolds equation. This approach of perturbed solutions has been discussed e.g. by Venner et. al. [16], Marusic et al. [10] and Tavakol et al. [15].
On substituting the ansatz Eqs. 15, 16 and 17 into Eqs. 4 and 5 at order $\varepsilon^{2}$, the velocity and pressure satisfy the following reduced equations,
(18) $\frac{\partial P_{2}}{\partial\bar{x}}$ $=\frac{\partial^{2}U_{0}}{\partial\bar{x}^{2}}+\frac{\partial^{2}U_{2}}{\partial\bar{y}^{2}},$
(19) $\frac{\partial P_{2}}{\partial\bar{y}}$ $=\frac{\partial^{2}V_{0}}{\partial\bar{y}^{2}}.$
@@
(20) $\frac{\partial P_{4}}{\partial\bar{x}}$ $=\frac{\partial^{2}U_{2}}{\partial\bar{x}^{2}}+\frac{\partial^{2}U_{4}}{\partial\bar{y}^{2}},$
(21) $\frac{\partial P_{4}}{\partial\bar{y}}$ $=\frac{\partial^{2}V_{2}}{\partial\bar{y}^{2}}+\frac{\partial^{2}V_{0}}{\partial\bar{x}^{2}}.$
@@
We assume the boundary conditions Eqs. 8, 9 and 10. Then $U_{0}$ and $V_{0}$ are the dimensionless forms of Eqs. 11 and 12. and $P_{0}$ is the dimensionless form of $p_{\text{Re}}$ obtained numerically. We define the flux $\mathcal{Q}$ of the $\varepsilon^{k}$-PLT solution to be consistent with that of the order $\varepsilon^{0}$ solution. Then in general, the pressure drop $\Delta\mathcal{P}$ of the $\varepsilon^{k}$-PLT solution will deviate from that of the order $\varepsilon^{0}$ solution.
The method of perturbed solutions is appealing in that we obtain exact expressions for the velocities $U_{k}$ and $V_{k}$ involving only $y$, the height function $h(x)$ and its derivatives. The expressions for pressure $P_{k}$ do involve integrals requiring numerical integration for an arbitrary height function. The complete expressions for $U_{k}$, $V_{k}$ and $P_{k}$ at orders $\varepsilon^{2}$ and $\varepsilon^{4}$ are presented in the appendix Appendix C. Similar expressions up to various orders appear in [16, 10, 15]. Here we have provided corrected expressions for $V_{2}$ and $U_{4}$ in [15], and we include an expression for $V_{4}$ which is not otherwise explicitly written in [16, 10, 15].

### v2

@@ Abstract
Lubrication theory makes use of the assumptions of a long and thin fluid domain and a small scaled Reynolds number to formulate a linearized approximation to the Navier-Stokes equations. Extended lubrication theory aims to improve the model accuracy by relaxing these assumptions and including additional terms in the formulation. However, such models are sensitive to large surface gradients which lead the assumptions of the model to break down. In this paper, we present a formulation of extended lubrication theory, and compare our model with several existing models, along with the numerical solution to the Stokes equations. The error in pressure and velocity is characterized for a variety of fluid domain geometries. Our results indicate that the new solution is suitable for a wide range of geometries. The magnitude of surface variation and the length scale ratio are both important factors influencing the accuracy of the extended lubrication theory models.
@@ 1 Introduction
The classical lubrication assumptions of a long and thin fluid domain and a small scaled Reynolds number are used to justify neglecting the inertial velocity and cross-film pressure gradient terms in the Navier-Stokes equations. The resulting Reynolds equation is a dramatic simplification of the original governing equations. More recently, variations on the thin film assumption have given rise to higher-order approximations for lubrication theory [venner_stokes_2003, marusic-paloka_second_2012, tavakol_extended_2017, takeuchi_extended_2019, sialmas_newtonian_2024, banerjee_variational_2024]. These extended lubrication theory models retain certain simplifying assumptions as in classical lubrication theory, while nevertheless being applicable to a broader scope of geometries in the low Reynolds number limit.
@@
Previous research on the flows modeled by lubrication theory has highlighted the sensitivity of lubrication theory to variations of the fluid film height. In particular, in the limit of zero Reynolds number flows, discrepancies between lubrication theory and Stokes flow increase with both the magnitude and frequency of large gradients in the height [brown_applicability_1995, biswas_backward-facing_2004, dobrica_about_2009]. The pressure drop modeled by the Reynolds equation may be underestimated due to additional pressure losses at points of sudden expansion that are not accurately modeled [dobrica_reynolds_2005, dobrica_about_2009]. Furthermore, notable flow features of corner recirculation seen in the Stokes solution are not accurately captured with the Reynolds equation [armaly_experimental_1983, biswas_backward-facing_2004, shyu_numerical_2018].
@@
In this paper, we consider classical lubrication theory (the Reynolds equation), and more recent variations on lubrication theory which we refer to as extended lubrication theory [takeuchi_extended_2019] and/or perturbed lubrication theory [venner_stokes_2003, marusic-paloka_second_2012, tavakol_extended_2017, sialmas_newtonian_2024] in the case of models based on perturbation theory. We asses the performance of these models at low Reynolds numbers by comparison to the Stokes equations. We propose a formulation of extended lubrication theory that builds on the model of Takeuchi & Gu (2019) [takeuchi_extended_2019]; our approach modifies the treatment of the velocity boundary conditions and ensures incompressibility. The pressure and velocity solutions from each model are compared in a variety of geometries: a logistic step, a backward facing step, and two versions of a triangular slider that we refer to as positive and negative based on the texture to be defined later on. Through these examples, we assess the sensitivity of each model to varying magnitudes of surface variation and discontinuities in the surface gradient in the low Reynolds number limit.
@@
Our results indicate that the extended and perturbed lubrication theory models provide significant improvements on classical lubrication theory in the case of small to moderate surface variation. For the logistic step, our proposed solution in extended lubrication theory has the smallest error in velocity for small slopes, and the smallest error in pressure at moderate slopes. For the triangular slider, the perturbed solutions have significantly reduced error in pressure, particularly in the case of negative texturing. In general, we find that for geometries with a large surface gradient, the extended and perturbed solutions may over-correct the cross-film pressure gradient.
@@ 2.1 Lubrication Theory
We begin with a review of classical lubrication theory. This treatment may be found in various textbooks [leal_advanced_2007], and we include it here to establish notation and to make the presentation self-contained. The Navier-Stokes equations,
@@
serve as the governing equations for a two-dimensional incompressible fluid. The pressure is $p$, the velocity is $(u,v)$, and the kinematic viscosity is $\nu=\eta/\rho$, where $\eta$ is the bulk viscosity and $\rho$ is the constant density.
Lubrication theory provides a linearized approximation to the Navier-Stokes equations valid for geometries in which the fluid is constrained to a relatively thin domain. For the two-dimensional fluid domain $(x,y)\in[x_{0},x_{f}]\times[0,h(x)]$, the characteristic length scales are $L_{x}=x_{f}-x_{0}$ and $L_{y}=\max h(x)>0$. Given a prescribed constant flux $\mathcal{Q}\neq 0$, the characteristic velocities are $U_{*}=\mathcal{Q}/L_{y}$ and $V_{*}=\mathcal{Q}/L_{x}$, and the characteristic pressure is $P_{*}=\eta U_{*}L_{x}/L_{y}^{2}$. The dimensionless variables are $X=x/L_{x}$, $Y=y/L_{y}$, $H=h/L_{y}$, $U=u/U_{*}$, $V=v/V_{*}$, and $P=p/P_{*}$. The Reynolds number is $\text{Re}=\rho U_{*}{\color[rgb]{0,0,0}L_{x}}/\eta$ and the length scale ratio is $\varepsilon=L_{y}/L_{x}$. The Navier-Stokes equations Eqs. 1 and 2 in dimensionless terms are,
(4) $\frac{\partial P}{\partial X}$ $=\Big(\varepsilon^{2}\frac{\partial^{2}U}{\partial X^{2}}+\frac{\partial^{2}U}{\partial Y^{2}}\Big)-\varepsilon^{2}\text{Re}\Big(U\frac{\partial U}{\partial X}+V\frac{\partial U}{\partial Y}\Big),$
(5) $\frac{\partial P}{\partial Y}$ $=\varepsilon^{2}\Big(\varepsilon^{2}\frac{\partial^{2}V}{\partial X^{2}}+\frac{\partial^{2}V}{\partial Y^{2}}\Big)-\varepsilon^{4}\text{Re}\Big(U\frac{\partial V}{\partial X}+V\frac{\partial V}{\partial Y}\Big),$
@@
(6) $U(X,0)=U_{b}$ $U(X,H)=0,$
(7) $V(X,0)=0$ $V(X,H)=0,$
where $U_{b}=\mathcal{U}/U_{*}$ is the constant relative velocity between the surfaces. We prescribe the flux $\mathcal{Q}$ so that, in dimensionless terms, the velocity satisfies,
(8) $\int_{0}^{H(X)}U(X,Y)\,dY=1.$
We also prescribe the fixed outlet pressure $P(1,0)=0$. Because the flux is prescribed, the average pressure drop $\Delta P$ is given from the solution, as in [banerjee_variational_2024],
(9) $\Delta P=\frac{1}{H(0)}\int_{0}^{H(0)}P(0,Y)\,dY-\frac{1}{H(1)}\int_{0}^{H(1)}P(1,Y)\,dY.$
Conversely, one may prescribe the average pressure drop, in which case the flux would be determined by the solution.
@@
(10) $\frac{\partial P_{0}}{\partial X}$ $=\frac{\partial^{2}U_{0}}{\partial Y^{2}},$
(11) $\frac{\partial P_{0}}{\partial Y}$ $=0,$
where $P_{0}$, $U_{0}$, $V_{0}$ denote the order $\varepsilon^{0}$ lubrication approximation. Together with incompressibility Eq. 3, Eqs. 10 and 11 constitute the governing equations for lubrication theory.
The velocity $U_{0}$ is determined through integration of Eq. 10 and applying the boundary conditions Eq. 6,
(12) $U_{0}(X,Y)=\frac{1}{2}\frac{dP_{0}}{dX}\Big(Y^{2}-HY\Big)+U_{b}\frac{\big(H-Y\big)}{H}.$
The velocity $V_{0}$ is determined from Eq. 12 and incompressibility Eq. 3 with the boundary conditions Eq. 7,
(13) $V_{0}(X,Y)=\frac{-1}{6}\frac{d^{2}P_{0}}{dX^{2}}Y^{3}+\frac{1}{2}\Bigg(\frac{1}{2}\Big(\frac{d^{2}P_{0}}{dX^{2}}H+\frac{dP_{0}}{dX}\frac{dH}{dX}\Big)-\frac{U_{b}}{H^{2}}\frac{dH}{dX}\Bigg)Y^{2}.$
The condition of constant flux and the boundary condition $V_{0}(X,H)=0$ are satisfied exactly when $P_{0}$ satisfies the Reynolds equation,
(14) $\frac{d}{dX}\Bigg[H^{3}\frac{dP_{0}}{dX}\Bigg]=6U_{b}\frac{dH}{dX}.$
To solve the Reynolds equation, we assign the inlet pressure gradient according to the prescribed flux, along with the fixed outlet pressure,
(15) $\frac{dP_{0}}{dX}\bigg|_{0}=-12\big[H(0)\big]^{-3}+6U_{b}\big[H(0)\big]^{-2},$ $P_{0}(1)=0.$
Note that the flux appears in the scaling of the boundary velocity, $U_{b}=\mathcal{U}L_{y}/\mathcal{Q}$. Since the flux is prescribed, the inlet pressure $P_{0}(0)$ and the average pressure drop $\Delta P_{0}$ are determined by the solution. The solution $P_{0}$ to the Reynolds equation is determined using a second-order accurate finite difference scheme, see Appendix A.
@@ 2.2 Perturbed Lubrication Theory
The dimensionless Navier-Stokes equations Eqs. 4 and 5 motivate the approach of an $\varepsilon$-perturbed solution [venner_stokes_2003, marusic-paloka_second_2012, tavakol_extended_2017, sialmas_newtonian_2024]. Given a sufficiently small Reynolds number such that $\varepsilon^{2}\text{Re}\ll 1$, perturbed lubrication theory assumes a solution of pressure and velocity of the form,
(16) $P(X,Y)$ $=P_{0}+\varepsilon^{2}P_{2}+\varepsilon^{4}P_{4},$
(17) $U(X,Y)$ $=U_{0}+\varepsilon^{2}U_{2}+\varepsilon^{4}U_{4},$
(18) $V(X,Y)$ $=V_{0}+\varepsilon^{2}V_{2}+\varepsilon^{4}V_{4}.$
The order $\varepsilon^{k}$ solution ($\varepsilon^{k}$-PLT) assumes a long and thin domain characterized by $\varepsilon^{k}\sim\mathcal{O}(1)$ and such that higher powers of $\varepsilon$ are negligible. The $\varepsilon^{0}$-PLT solution $U_{0}$, $V_{0}$, $P_{0}$ is the solution of lubrication theory and the Reynolds equation.
On substituting the ansatz Eqs. 16, 17 and 18 into Eqs. 4 and 5 at order $\varepsilon^{2}$, the velocity and pressure satisfy the following reduced equations,
(19) $\frac{\partial P_{2}}{\partial X}$ $=\frac{\partial^{2}U_{0}}{\partial X^{2}}+\frac{\partial^{2}U_{2}}{\partial Y^{2}},$
(20) $\frac{\partial P_{2}}{\partial Y}$ $=\frac{\partial^{2}V_{0}}{\partial Y^{2}}.$
@@
(21) $\frac{\partial P_{4}}{\partial X}$ $=\frac{\partial^{2}U_{2}}{\partial X^{2}}+\frac{\partial^{2}U_{4}}{\partial Y^{2}},$
(22) $\frac{\partial P_{4}}{\partial Y}$ $=\frac{\partial^{2}V_{2}}{\partial Y^{2}}+\frac{\partial^{2}V_{0}}{\partial X^{2}}.$
@@
The order $\varepsilon^{0}$ Reynolds solution $U_{0}$, $V_{0}$, and $P_{0}$ satisfies the boundary conditions Eqs. 6, 7 and 15 respectively. The higher-order velocities $U$ and $V$ also satisfy Eqs. 6 and 7. The higher-order pressures $P$ satisfy $P(1,0)=0$, but do not necessarily satisfy the pressure gradient condition in Eq. 15. Instead, we require that the flux $\mathcal{Q}$ be the same as the flux of the order $\varepsilon^{0}$ solution, such that, in dimensionless terms, $\int_{0}^{H}U_{0}\,dY=1$ and $\int_{0}^{H}U_{k>0}\,dY=0$. Because the flux is prescribed, the average pressure drop $\Delta P_{\varepsilon^{k}\text{-PLT}}$ is determined by the solution. In general, the average pressure drops $\Delta P_{\varepsilon^{2}\text{-PLT}}=\Delta P_{0}+\varepsilon^{2}\Delta P_{2}$ and $\Delta P_{\varepsilon^{4}\text{-PLT}}=\Delta P_{0}+\varepsilon^{2}\Delta P_{2}+\varepsilon^{4}\Delta P_{4}$ will deviate from $\Delta P_{0}$. We determine $\Delta P_{k>0}$ through numerical integration according to Eq. 9.
The method of perturbed solutions is appealing in that we obtain exact expressions for the velocities $U_{k}$ and $V_{k}$ involving only $Y$, the height function $H$ and its derivatives. The expressions for pressure $P_{k}$ do involve integrals requiring numerical integration for an arbitrary height function. The complete expressions for $U_{k}$, $V_{k}$ and $P_{k}$ at orders $\varepsilon^{2}$ and $\varepsilon^{4}$ are presented in the appendix Appendix C. Similar expressions up to various orders appear in [venner_stokes_2003, marusic-paloka_second_2012, tavakol_extended_2017]. Here we have provided corrected expressions for $V_{2}$ and $U_{4}$ in [tavakol_extended_2017], and we include an expression for $V_{4}$ which is not otherwise explicitly written in [venner_stokes_2003, marusic-paloka_second_2012, tavakol_extended_2017].
The method of perturbed solutions is also well suited to techniques such as diagonal Padé approximants or the Shanks’ transform [sialmas_newtonian_2024]. These techniques may be used to accelerate and extend the domain of the convergence for the series solution. Although such numerical methods are out of the scope of the present work, we consider this a potential avenue for future investigation and improvement of the PLT method.